OpenStudy (anonymous):

An airplane flies horizontally at a constant velocity vx = 115m/s and height h = 1050m. It drops a care package.  The magnitude and direction of the velocity vector of the package when it reaches the ground is closest in value to: A 115m=s and 51 B 143m=s and 51 C 184m=s and 51 D 184m=s and 309

OpenStudy (anonymous):

1st use h=vi*t + .5at^2 to solve for t because a = -9.81m/s^2 and vi =0 then with t use V = Vi + at to find the y component of the Packages velcoity V = (Vx^2 + Vy^2)^(1/2) Angle = arctan (vy/vx)

OpenStudy (anonymous):

Thank you so much. I like your drawing too @demitris